40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 18:49:44 2013, Gautam, Update, General, ITMx Oplev Fixed  
    Reply  Wed Jun 5 20:45:19 2013, gautam, Update, General, ITMx Oplev Fixed  ITMx_overview.JPGITMx_clipping_.JPG
       Reply  Wed Jun 26 19:02:38 2013, gautam, Update, General, ITMx Oplev (not quite?) Fixed oplev_chamber.pdfphoto_1.JPGphoto_2.JPG
          Reply  Thu Jun 27 18:11:53 2013, gautam, Update, General, ITMx Oplev-POX looks beam okay 
             Reply  Thu Jun 27 23:01:39 2013, gautam, Update, General, ITMx Oplev-servo gains adjusted pitch-plot.pdfyaw-plot.pdf
                Reply  Fri Jun 28 13:10:28 2013, gautam, Update, General, ITMx Oplev-servo gains adjusted pitch-plot_copy.pdfyaw-plot_copy.pdf
Message ID: 8784     Entry time: Fri Jun 28 13:10:28 2013     In reply to: 8777
Author: gautam 
Type: Update 
Category: General 
Subject: ITMx Oplev-servo gains adjusted 

 

 As mentioned in elog 8770, I wanted to give the POX beam a little more clearance from the pick-off mirror steering the outcoming oplev beam. I tweaked the position of this mirror a little this morning, re-centred the spot, and checked the loop transfer function once again. These were really close to those I measured last night (UGF for pitch ~8Hz, for yaw ~7Hz), reported in elog 8777, so I did not have to change the loop gains for either pitch or yaw. Plots attached.

 pitch-plot_copy.pdf

 

yaw-plot_copy.pdf

ELOG V3.1.3-